Consider the given state of stress. Take X = 10 MPa and Y = 45 MPa.

60 MPa

Determine the principal planes using Mohr's circle. a) The principal planes are at −  ° and  °.

Determine the principal stresses using Mohr's circle. b)The minimum principal stress is −  MPa, and the maximum principal stress is  MPa.

Determine the orientation of the planes of maximum in-plane shearing stress using Mohr's circle. c) The orientation of the plane of maximum in-plane shearing stress in the first quadrant is  °. The orientation of the plane of maximum in-plane shearing stress in the second quadrant is  °.

Determine the maximum in-plane shearing stress using Mohr's circle. d) The maximum in-plane shearing stress is  MPa.

Determine the normal stress using Mohr's circle. e)The normal stress is  MPa.
Answer #1

Given State of stren Standard Element 45MPa Gorpa lorpa . Tx= LOMPa Ty= -45hPa ox, ty are & They tue Тау -- бонро- 97 (+) Th6 Major Ę Mindt principle Streu o 7 ox toy t Pay oz on 100k + sleeus )* + 1-4 10-45 To +45 12 -11: St | 21^+ 65 T =) - 17.5+.-120 Taul20p) = 55 2 Op =-65.37 - 32.688 qe oport Since principal Planes are Opz= -32.688 +90% [Opz = 57.312 t О - Нахімом іOs2= Op. +45° - 57.312 + 45° [Os2 = 10233120 and quadrant © Holmeal stren= Extry Average - lo-u5 = 2 To = -17.5 MPa